OpenStudy (anonymous):

In the number 0.148285, the digit 148285 repeated indefinitely. how many of the first 500 digits after the decimal point are odd? A) 83 B) 166 C) 167 D) 168 E) 332

myininaya (myininaya):

so 6 digits are repeated over and over and we want to look at the first 500 digits 500/6=83(6)+2 so we will see all those numbers over and over and at the very end we will see 14 (only one of these digits are odd) so looking at the first 498 digits we have that 2 of each of the 6 of the 83 are odd so we have 2*83+1=167